A person runs 800m at an average speed of 6.0m/s and then the next 800 M at an average speed of 8.0 m/s. What is the total time required?
1
Expert's answer
2020-06-11T10:44:44-0400

The total time required is the sum of distances divided by the corresponding speed:


"t=\frac{d_1}{s_1}+\frac{d_2}{s_2}=\frac{800}{6}+\frac{800}{8}=233\text{ s},"

or 3 min 53 s.